# What is Dataset Distillation Learning?

### v1

@@ 1 Introduction
The past decade in machine learning research has been marked with incredible breakthroughs in leveraging over-parameterized models. As a consequence, the landscape of machine learning research has been increasingly dominated by massive datasets. The growing size of training datasets presents a new challenge on the infrastructure required to store and train on such data. Large-scale data not only strains existing compute infrastructure with increased training times, but also limits its accessibility to researchers with sufficient compute infrastructure. Therefore, there is a crucial need for scaling down large-scale datasets.
Figure 1: Real vs. distilled data. Real images of airplane, car, and truck from CIFAR-10 (Krizhevsky et al., 2009) are shown on left and highly salient distilled images of the same classes are shown on the right. While distilled images can be used to train high-accuracy classifier, why this is possible and what do they represent remains unclear.
@@
To what extent can distilled data act as substitute for real data? Dataset distillation is commonly motivated by synthesizing information rich data that can serve as effective drop-in replacements of the original dataset (Sachdeva & McAuley, 2023). However, this comes with limitations and trade-offs. Prior works have already identified one such limitation: data distilled using one model architecture can not be effectively used to train a different model architecture (Zhong & Liu, 2023). We perform additional analysis to reveal the extent of distilled data’s ability to substitute for real data. We demonstrate that models trained on real data are able to successfully recognize the classes in distilled data, demonstrating that distilled data does encode transferable semantics. However, at training time, simply mixing real data with distilled data results in decrease in the performance of the final classifier. Therefore, distilled data should not be treated as real data during training, and we have to be careful in training with distilled data outside the typical evaluation setting for dataset distillation (training only on distilled data and on the same model architecture).
@@
Do distilled data points idividually carry meaningful information? Given that the whole distilled dataset compresses the early training dynamics, are individual examples still meaningful? To answer this question, we introduce a new interpretable framework for distilled data by leveraging a popular interpretability method: influence functions (Koh & Liang, 2017). In contrast to their intended uses to better understand a model, we utilize influence functions to better understand the data. We empirically demonstrate the power and consistency of the framework and reveal that notable semantic information is stored in different distilled data points: for example, one distilled image is associated with classifying yellow cars whereas another with cars in parking lots.
@@ 2 Preliminary
Distribution matching (2) also distill data at a class level and uses penultimate features from random convolutional neural networks as $\mathcal{F}^{p}$ and similarity function $D$ defined as the maximum mean discrepancy (Gretton et al., 2012) to arrive at the objective $\operatorname*{arg\,min}_{\mathcal{X}_{s}}D[\mathcal{F}^{p}(\mathcal{X}_{s}),\mathcal{F}^{p}(\mathcal{X}_{r})]$.
@@
We leverage the CIFAR-10 (Krizhevsky et al., 2009) dataset for our analysis with additional analysis on other datasets in Section B of the appendix. We use the standard three layers deep, 128 filters wide convolutional neural networks to train on distilled data and real data with 0.01 learning rate and 0.9 momentum for 300 iterations using SGD optimizer. We leverage the distilled data provided by the authors for gradient matching (3) and trajectory matching (4) while we reproduce the distilled data for BPTT (1) and distribution matching (2). In contrast to the original BPTT paper, we distill images initialized from real images rather than uniform Xavier initialization (Glorot & Bengio, 2010) since it gives better behaved distilled data (details provided in section F of the appendix).
@@ 5 Semantics of Captured Information
Figure 8: Influence of distilled images is consistent. left. Kernel density estimation plot of distribution of Pearson correlation between influences calculated across two different random seed. The high correlation confirms that the observed influence is due to distilled image rather than random chance. right. The correlation between the influence of a distlled image calculated with leave-one-out retraining when using 100 real images versus when using the 99 other distilled images. The moderately high correlations suggest that the influence of a distilled image extends to real data training, and hence, the information compressed is universal and not contingent on other distilled images.
@@ 5.1 Influence functions for understanding distilled data
To confirm the necessity of calculating the influence for examination of information content, in Figure 7, we verify that influence functions reveal information that cannot be glimpsed through visual similarity alone. Although real images $x_{t}$ which are identified as being highly influential do tend to be somewhat visually similar to the corresponding distilled image $x_{d}$ (Figure 7 bottom right), overall there is only a very weak correlation between the two measures. Looking at a visual example of a distilled image along with two highly-ranked real images (by closest distance and highest influence; Figure 7 top), we begin to notice that highest influence images do appear to be semantically consistent (e.g., both are images of orange cars). In Sections 5.2 we leverage this insight to begin to understand the information content of distilled images.
@@ 6 Conclusion
Through our investigation, we have addressed three key questions. First, we showed that while distilled data behave like real data at inference time, they are sensitive to the training procedure and cannot serve as drop in replacements for real data. Second, we demonstrated that dataset distillation captures the early learning dynamics of real models. Finally, we revealed that individual distilled data points encapsulate meaningful semantic information. Our study offers valuable insights into underlying mechanism of dataset distillation, informing better design of future methods.
@@ Appendix B Additional Mixing and Prediction Analysis
In this section, we detail additional analysis on the mixing experiment done in Figure 3 right as well prediction analysis done in Figure 4 to show our findings generalizes to different datasets, larger scales, and state-of-the-art (SOTA) methods.

### v2

@@ 1 Introduction
The past decade in machine learning research has been marked with incredible breakthroughs in leveraging over-parameterized models. As a consequence, the landscape of machine learning research has been increasingly dominated by massive datasets. The growing size of training datasets presents a new challenge on the infrastructure required to store and train on such data. Large-scale data not only strains existing compute infrastructure with increased training times but also limits its accessibility to researchers with sufficient compute infrastructure. Therefore, there is a crucial need for scaling down large-scale datasets.
Figure 1: Real vs. distilled data. Real images of airplane, car, and truck from CIFAR-10 (Krizhevsky et al., 2009) are shown on left and highly salient distilled images of the same classes are shown on the right. While distilled images can be used to train high-accuracy classifiers, why this is possible and what do they represent remains unclear.
@@
To what extent can distilled data act as substitute for real data? Dataset distillation is commonly motivated by synthesizing information rich data that can serve as effective drop-in replacements of the original dataset (Sachdeva & McAuley, 2023). However, this comes with limitations and trade-offs. Prior works have already identified one such limitation: data distilled using one model architecture can not be effectively used to train a different model architecture (Zhong & Liu, 2023). We perform additional analyses to reveal the extent of distilled data’s ability to substitute for real data. We demonstrate that models trained on real data are able to successfully recognize the classes in distilled data, demonstrating that distilled data does encode transferable semantics. However, at training time, simply mixing real data with distilled data results in decrease in the performance of the final classifier. Therefore, distilled data should not be treated as real data during training, and we have to be careful in training with distilled data outside the typical evaluation setting for dataset distillation (training only on distilled data and on the same model architecture).
@@
Do distilled data points individually carry meaningful information? Given that the whole distilled dataset compresses the early training dynamics, are individual examples still meaningful? To answer this question, we introduce a new interpretable framework for distilled data by leveraging a popular interpretability method: influence functions (Koh & Liang, 2017). In contrast to their intended uses to better understand a model, we utilize influence functions to better understand the data. We empirically demonstrate the power and consistency of the framework and reveal that notable semantic information is stored in different distilled data points: for example, one distilled image is associated with classifying yellow cars whereas another with cars in parking lots.
@@ 2 Preliminary
Distribution matching (2) also distills data at a class level and uses penultimate features from random convolutional neural networks as $\mathcal{F}^{p}$ and similarity function $D$ defined as the maximum mean discrepancy (Gretton et al., 2012) to arrive at the objective $\operatorname*{arg\,min}_{\mathcal{X}_{s}}D[\mathcal{F}^{p}(\mathcal{X}_{s}),\mathcal{F}^{p}(\mathcal{X}_{r})]$.
@@
We leverage the CIFAR-10 (Krizhevsky et al., 2009) dataset for our analysis with additional analyses on other datasets in Section B of the appendix. We use the standard three layers deep, 128 filters wide convolutional neural networks to train on distilled data and real data with 0.01 learning rate and 0.9 momentum for 300 iterations using SGD optimizer. We leverage the distilled data provided by the authors for gradient matching (3) and trajectory matching (4) while we reproduce the distilled data for BPTT (1) and distribution matching (2). In contrast to the original BPTT paper, we distill images initialized from real images rather than uniform Xavier initialization (Glorot & Bengio, 2010) since it gives better behaved distilled data (details provided in section F of the appendix).
@@ 5 Semantics of Captured Information
Figure 8: Influence of distilled images is consistent. left. Kernel density estimation plot of distribution of Pearson correlation between influences calculated across two different random seeds. The high correlation confirms that the observed influence is due to distilled image rather than random chance. right. The correlation between the influence of a distlled image calculated with leave-one-out retraining when using 100 real images versus when using the 99 other distilled images. The moderately high correlations suggest that the influence of a distilled image extends to real data training, and hence, the information compressed is universal and not contingent on other distilled images.
@@ 5.1 Influence functions for understanding distilled data
To confirm the necessity of calculating the influence for examination of information content, in Figure 7, we verify that influence functions reveal information that cannot be glimpsed through visual similarity alone. Although real images $x_{t}$ which are identified as being highly influential do tend to be somewhat visually similar to the corresponding distilled image $x_{d}$ (Figure 7 bottom right), overall there is only a very weak correlation between the two measures.
Looking at a visual example of a distilled image along with two highly-ranked real images (by closest distance and highest influence; Figure 7 top), we begin to notice that highest influence images do appear to be semantically consistent (e.g., both are images of orange cars). In Sections 5.2 we leverage this insight to begin to understand the information content of distilled images.
@@ 6 Conclusion
Through our investigation, we have addressed three key questions. First, we showed that while distilled data behave like real data at inference time, they are sensitive to the training procedure and cannot serve as drop in replacements for real data. Second, we demonstrated that dataset distillation captures the early learning dynamics of real models. Finally, we revealed that individual distilled data points encapsulate meaningful semantic information. Our study offers valuable insights into underlying mechanisms of dataset distillation, informing better design of future methods.
@@ Appendix B Additional Mixing and Prediction Analysis
In this section, we detail additional analyses on the mixing experiment done in Figure 3 right as well prediction analysis done in Figure 4 to show our findings generalizes to different datasets, larger scales, and state-of-the-art (SOTA) methods.